mudasali69:

A man spent 10% of his money. After spending 60% of the remainder he has Rs.72 left. How much had he in the start?

blitzo5000:

142 10+60=70+72+142

blitzo5000:

=142 srry

Vocaloid:

You cannot add the number of rupees to the percentage of rupees. They represent different things. Let’s suppose the original number of rupees is x. He spent 10% of the original amount, so he spent 0.10x and thus has 1x - 0.1x = 0.9x left. He spent 60% of the remainder, meaning he kept 100% - 60% = 40% of the remainder so he has (0.40)(0.9x) = 72 rupees left. Solve for x.

blitzo5000:

bruh what ever